OpenStudy (theviper):

Please Help:) The value of 'k' for which "x+k" is a factor of\[\LARGE{\color{blue}{x^3+kx^2-2x+k+4}}\]is???? @TuringTest @cshalvey @KingGeorge @nbouscal @dumbcow @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

If x+k is a factor of x^3+kx^2-2x+k+4 Then x = -k is a zero of x^3+kx^2-2x+k+4 So replace x with -k and the result should be zero So x^3+kx^2-2x+k+4 (-k)^3+k(-k)^2-2(-k)+k+4 = 0 Solve that for k

OpenStudy (theviper):

\[\Huge{\color{green}{k=\frac{-4}{3}}}\]???? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (theviper):

thanx:)

jimthompson5910 (jim_thompson5910):

yw